Alcohol consumption slows peoples reaction times. In a controlled government test, it takes a certain driver t - 0.35 s to hit the brakes in a crisis when unimpaired and 12 1.05 s when drunk. If the car is initially traveling at v 87 km/h, how much farther does the car travel before the brakes are applied when the person is drunk than it travels when the person is sober? additional distance traveled TOOLS x10
Answer #1

Initial speed, v=87km/h=24.17m/s

final speed, v_{f}=0m/s

Unimpaired driver

ti 0.35s

Use Formula v=u+at

v_{f}=v+a_{1}t_{1}

0m/s= 24.17m/s+ al * 0.35s

a69.06m/s2

Use Formula sut +1/2at2

s_{1}=vt_{1}+1/2a_{1}t_{1}^{2}

Sj = 24.17m/s * 0.35s-1/2 * 69.06m/s2 * (0.35s)2

s_{1}=24.17*0.35-0.5*69.06*0.35^{2}

{color{Red} s_{1}=4.23m}

------------------------------------

Drunk driver

t_{2}=1.05s

Use Formula v=u+at

v_{f}=v+a_{2}t_{2}

Om/s-24.17m/s + a2 *1.05s

23.02m/s

Use Formula sut +1/2at2

s_{2}=vt_{2}+1/2a_{2}t_{2}^{2}

s_{2}=24.17m/s*1.05s-1/2*23.02m/s^{2}*(1.05s)^{2}

s_{2}=24.17*1.05-0.5*23.02*1.05^{2}

{color{Red} s_{2}=12.69m}

------------------------------------

s=s_{2}-s_{1}=12.69m-4.23m

ANSWER: s 8.4677 |

====================